200 grams metal at 1000C is placed on calorimeter which contains 300 grams water at 180C. If the specific heat of metal is 140 J/kg0C, determine the final temperature of the system!

1
Expert's answer
2022-04-04T09:03:38-0400

Given:

"m_1=200\:\rm g"

"t_1=100^\circ\rm C"

"m_2=300\:\rm g"

"t_2=18^\circ\rm C"

"c_1=140\:\rm J/kg^\circ\rm C"

"c_2=4184\:\rm J/kg^\circ\rm C"

The heat balance equation says

"c_1m_1(t_1-t_3)=c_2m_2(t_3-t_2)"

Hence, the final temperature of system

"t_3=\frac{c_1m_1t_1+c_2m_2t_2}{c_1m_1+c_2m_2}"

"t_3=\frac{140*0.2*100+4184*0.3*18}{140*0.2+4184*0.3}=20^\circ\rm C"